asierra599:

Elena agrees to finish a knitting project for a friend. The graph shows the number of rows Elena completes compared to the amount of time spent knitting. Assuming she worked at a constant rate, how many rows had been completed before Elena started working? 12 14 15 19

Vocaloid:



Vocaloid:

@dude would you mind taking a look at this when you get a chance

dude:

Well, you have to find the equation of the line for the points given [They told you this was a linear function as they mentioned it was at a constant rate] y=mx+b To find m, use \(\dfrac{y_2-y_1}{x_2-x_1}\) [Choose any two points] I would recommend using (14, 19) ->(\(x_1,y_1\)) and (20, 22) ->(\(x_2,y_2\)) \(\dfrac{22-19}{20-14}=\dfrac{3}{6}=\dfrac{1}{2}\) We have \(y=\dfrac12x+b\) Now, we can substitute a point they gave us into x and y to solve for b (which is the value we really want) [I used (20,22) here but you can use any other point] \(22=\dfrac12(20)+b\) Solve for b and that should be your starting amount b = y-intercept = starting point